# Extending class group action attacks via sesquilinear pairings

### v1

@@ 2.3 Computational assumptions
With regards to computations, we use the word efficient to mean polynomial time in the size of the input. Throughout the paper, when we assume that if we are given an $\mathcal{O}$-oriented elliptic curve, we mean that we are given an explicit orientation, and that, in particular that, given an element $\alpha\in\mathcal{O}$, we can compute its action $[\alpha]$ on a point $P$ on $E$ efficiently.
@@
We assume that $m$ is smooth, meaning that its factors are polynomial in size, so that discrete logarithms in $\mu_{m}$ or $E[m]$ are computable in polynomial time. In particular, we can write any element of $E[m]$ in terms of a given basis.
@@
Let $m>1$ be an integer. Let $E$ be an elliptic curve defined over a field $\mathbb{F}$ (assumed finite in this paper). Suppose that $P\in E(\mathbb{F})[m]$. Choose divisors $D_{P}$ and $D_{Q}$ of disjoint support such that $D_{P}\sim(P)-(\mathcal{O})$ and $D_{Q}\sim(Q)-(\mathcal{O})$. Then $mD_{P}\sim 0$, hence there is a function $f_{P}$ such that $\operatorname{div}(f_{P})=mD_{P}$. The Tate-Lichtenbaum pairing
@@
The standard properties of the Tate pairing are as follows. Proofs can be found in many places, for example [22] and [4, Sec 3.2].
@@ 3 Structure of $E[\alpha]$
Suppose first that $\alpha$ is separable. Let $\iota$ be an $\mathcal{O}$-orientation for $E$ and $\mathcal{O^{\prime}}$ be the order of $K$ for which $\iota$ is a primitive orientation. Then as an abelian group the quotient $\operatorname{End}(E)/\mathcal{O^{\prime}}$ is torsion-free and Theorem 3.1 tells us that $E[\alpha]\cong\mathcal{O^{\prime}}/\alpha\mathcal{O^{\prime}}$ as $\mathcal{O}^{\prime}$-modules. We have $N(\alpha)=N(\alpha\mathcal{O^{\prime}})=\left|\mathcal{O^{\prime}}/\alpha\mathcal{O^{\prime}}\right|$, so since $N(\alpha)$ is coprime to $f$, it follows from [11, Proposition 7.18, 7.20] that the natural injection $\mathcal{O}/\alpha\mathcal{O}\rightarrow\mathcal{O^{\prime}}/\alpha\mathcal{O^{\prime}}$ is an isomorphism of $\mathcal{O}$-modules.
@@
Finally, suppose $\alpha$ factors through $[n]$ for some $n>1$ with $n\mid f$. Then as a $\mathbb{Z}$-module, $E[\alpha]\cong\mathbb{Z}/b\mathbb{Z}\times\mathbb{Z}/c\mathbb{Z}$ with $n\mid b\mid c$. Let $\{P,Q\}$ be a generating set for $E[\alpha]$ with $\operatorname{ord}(P)=b,\operatorname{ord}(Q)=c$ and let $\mathcal{O^{\prime}}=\mathbb{Z}[\sigma]$ for some $\sigma$. Then $\mathcal{O}=\mathbb{Z}[f\sigma]$. Since any element of $\mathcal{O}$ is a $\mathbb{Z}$-linear combination of $[1]$ and $[f\sigma]$, whether or not $E[\alpha]$ is cyclic as an $\mathcal{O^{\prime}}$-module is determined by the action of $f\sigma$. We have $[f\sigma]Q=[n\sigma]Q^{\prime}$, where $Q^{\prime}=[f/n]Q$.
If $[\sigma]Q^{\prime}=[s]P+[t]Q$, then $[f\sigma]Q=[ns]P+[nt]Q$, and we cannot obtain $P$ from the action of any $\mathbb{Z}$-linear combination of $[1]$ and $[f\sigma]$ on $Q$. Thus, $Q$ cannot be a generator for $E[\alpha]$ as an $\mathcal{O}$-module. Since $Q$ was an arbitrary order $c$ point, and since no point of order strictly less than $c$ can generate $E[\alpha]$ as an $\mathcal{O}$-module (endomorphisms send points of $E$ of order $m$ to points of $E$ of order dividing $m$), $E[\alpha]$ cannot be a cyclic $\mathcal{O}$-module.
@@
Consider the ordinary curve $y^{2}=x^{3}+30x+2$ over $\mathbb{F}_{101}$. Denoting the Frobenius endomorphism by $\pi$, $\mathbb{Z}[\pi]$ has conductor 2 in the maximal order and $[\mathbb{Z}[\pi]:\mathbb{Z}[\pi^{2}]]=18$. Thus, Theorem 3.3 implies $E[3]$ is not cyclic as a $\mathbb{Z}[\pi^{2}]$-module. Indeed, making a base change to $\mathbb{F}_{101^{2}}$, the 3-torsion of $E$ is rational. On the other hand, $E[3]$ is cyclic as a $\mathbb{Z}[\pi]$-module. With $\mathbb{F}_{101^{2}}=\mathbb{F}_{101}(a)$ and $x^{2}-4x+2$ the minimal polynomial of $a$, we have $P=(41a+16,39a+19)\in E[3]$ and $\pi(P)=(60a+79,62a+74)\not\in\langle P\rangle$, hence $\mathbb{Z}[\pi]P=E[3]$.
@@ 4 Sesquilinear pairings
We follow [26] in this section. Suppose $\mathcal{O}=\mathbb{Z}[\tau]$ is an imaginary quadratic order. Let $E$ have CM by $\mathcal{O}$. Let $\rho:\mathcal{O}\rightarrow\operatorname{GL}_{2}(\mathbb{Z})$ be the left-regular representation of $\mathcal{O}$ acting on the basis $1$ and $\tau$. Then we endow the Cartesian square $(\mathbb{F}^{*})^{\times 2}$ with an $\mathcal{O}$-module action via
$(x,y)^{\alpha}=\rho(\alpha)\cdot(x,y)=(x^{a}y^{b},x^{c}y^{d}),\quad\text{where}\quad\rho(\alpha)=\begin{pmatrix}a&b\\ c&d\end{pmatrix}.$
@@
$\operatorname{div}(f_{P,1})$ $=a([-\tau]P)+b(P)-(a+b)(\mathcal{O}),$
$\operatorname{div}(f_{P,2})$ $=c([-\tau]P)+d(P)-(c+d)(\mathcal{O}).$
Choose an auxiliary point $R$ and define
@@
Then, choosing $R$ so that the necessary supports are disjoint (i.e. the support of $\operatorname{div}(f_{P,i})$ and $D_{Q,j}$ are disjoint for each pair $i$, $j$), the pairing is defined as
@@
In [26], it is shown how it is possible to think of these definitions as divisors in $\mathcal{O}\otimes_{\mathbb{Z}}\operatorname{Pic}^{0}(E)$:
$D_{Q}=D_{Q,1}+\tau\cdot D_{Q,2},\quad D_{P}=([-\tau]P)-(\mathcal{O})+\tau\cdot((P)-(\mathcal{O})),$
and analogously define $f_{P}$ satisfying $\operatorname{div}(f_{P})={\alpha}\cdot D_{P}$, so that the definition above has the form $f_{P}(D_{Q})$ as for the classical Tate pairing. For simplicity here, we stick to the direct definition above. In that same paper, analogous constructions are also given for quaternion orders and Weil-like pairings.
Theorem 4.1 ([26, Theorems 5.4, 5.5, 5.6])
@@
    $\widehat{T}^{\tau}_{N(\alpha)}(P,Q)=\widehat{T}^{\tau}_{\alpha}(P,Q)^{\overline{\alpha}}\pmod{((\mathbb{F}^{*})^{\times r})^{\alpha}}$
@@
The pairing $\widehat{T}_{\alpha}^{\tau}(P,Q)$ is efficiently computable. That is, it takes polynomially many operations in the field of definition of $P$ and $Q$.
@@
This follows from the definition given above, which is amenable to a Miller-style pairing algorithm; details are in [26, Algorithm 5.7].
@@
Consider the image $N(\mathcal{O})$ of $\mathcal{O}$ under the norm map. Then $N(\mathcal{O})$ modulo $m$ is a subset of $\{x^{2}:x\in\mathbb{Z}/m\mathbb{Z}\}$ only if $m$ and $\Delta_{\mathcal{O}}$ share a non-trivial factor.
@@
We may assume, by Sunzi’s Theorem (Chinese Remainder Theorem), that $m$ is a prime power $p^{k}$. If $p$ is split, the statement follows from the fact that the norm map from $\mathbb{Q}_{p}\otimes_{\mathbb{Z}}\mathcal{O}$ to $\mathbb{Q}_{p}$ is surjective. If $p$ is inert, the norm map from $\mathbb{Q}_{p}\otimes_{\mathbb{Z}}\mathcal{O}$ to $\mathbb{Q}_{p}$ is not surjective, but is surjective on the residue field, so $N(\mathcal{O})$ modulo $p^{k}$ does include non-squares.
@@
Let $E$ be an elliptic curve oriented by $\mathcal{O}=\mathbb{Z}[\tau]$. Let $m$ be coprime to the discriminant $\Delta_{\mathcal{O}}$. Let $\mathbb{F}$ be a finite field containing the $m$-th roots of unity. Suppose $E[m]=E[m](\mathbb{F})$. Let $P$ have order $m$. Let $s$ be the maximal divisor of $m$ such that $E[s]\subseteq\mathcal{O}P$. Then the multiplicative order $m^{\prime}$ of $\widehat{T}_{m}^{\tau}(P,P)$ satisfies $s\mid m^{\prime}\mid s^{2}$. In particular, if $\mathcal{O}P=E[m]$, then the self-pairing has order $m$; and if $\mathcal{O}P=\mathbb{Z}P$, then the self-pairing is trivial.
@@
Let $m^{\prime}\mid m$ be the order of $\widehat{T}_{m}^{\tau}(P,P)$. Suppose $s$ is the maximal divisor of $m$ so that $E[s]\subseteq\mathcal{O}P$. Then, in particular, $[s\tau]P\in\mathbb{Z}P$. Thus $\mathcal{O}[s]P=\mathbb{Z}[s]P$. We will show that $m^{\prime}\mid s^{2}$ and $s\mid m^{\prime}$.
Let $\lambda\in\mathcal{O}$. Then $[\lambda s]P=[ks]P$ for some $k=k(\lambda)\in\mathbb{Z}$, and then
@@
Ranging over all $\lambda\in\mathcal{O}$, we conclude that $N(\lambda)$ are squares modulo $m^{\prime\prime}=m^{\prime}/\gcd(m^{\prime},s^{2})$, the multiplicative order of $\widehat{T}_{m}^{\tau}([s]P,[s]P)$, contradicting that $m$ is coprime to the discriminant unless $m^{\prime\prime}=1$ by Lemma 1. Therefore $m^{\prime}\mid s^{2}$.
@@
This has order $s$ on the left. Therefore $\widehat{T}_{m}^{\tau}(P,P)$ must have order a multiple of $s$. Hence $s\mid m^{\prime}$.
@@
In [4], the authors study cyclic self-pairings on oriented elliptic curves, namely homogeneous functions $f:\langle P\rangle\rightarrow\overline{\mathbb{F}}^{*}$ of degree $2$ which are compatible with endomorphisms from the orientation, in the sense that $\langle P\rangle$ is stabilized by such endomorphisms and $f(\phi P)=f(P)^{\deg\phi}$ for such endomorphisms. They show that non-trivial such pairings can only exist for $P$ of order $m$ dividing $\Delta_{\mathcal{O}}$. The reason our pairings are not ruled out by this result is that our pairings need not be defined on a cyclic subgroup stabilized by the orientation. In fact, their result can be recovered from Theorem 4.3, which implies that the pairings $\widehat{T}_{\alpha}^{\tau}$ are trivial on cyclic subgroups stabilized by $\mathcal{O}$.
@@
Let $E$ be an elliptic curve defined over a finite field $\mathbb{F}$, and let $m\in\mathbb{Z}$. Let a basis for $E[m]$ be given. Suppose arithmetic in $\mathbb{F}$, discrete logarithms in $\mathbb{F}^{*}$ modulo $m$, and group law computations on $E[m]$ can all be accomplished in polynomial time. Suppose $E$ is known to be oriented by $\mathcal{O}=\mathbb{Z}[\tau]$ (but the orientation $\iota$ is not given), and suppose $m$ is coprime to the discriminant $\Delta_{\mathcal{O}}$. Then the computation of arbitrary pairings $\widehat{T}^{\tau}_{m}(P,Q)$ on $E[m]$ is equivalent in polynomial time to the computation of the action of $[\tau]$ on $E[m]$.
@@
Note that since $\tau$ is known abstractly, computation of $[\tau]$ allows for computation of $[\overline{\tau}]=[Tr(\tau)]-[\tau]$.
@@
Conversely, suppose one can compute $\widehat{T}^{\tau}_{m}(P,Q)$ for any $P,Q\in E[m]$. We will show how to compute the action of $[\tau]$ on $E[m]$. The pairing is non-degenerate as a consequence of the given hypotheses. Choose $P\in E[m]$ uniformly randomly as a linear combination of the given basis. Compute $\widehat{T}^{\tau}_{m}(P,P)$. If it is not of order $m$, choose another $P$.
@@
If $\mathcal{O}P\neq E[m]$, then the algorithm is not guaranteed to be correct. Therefore, we run the algorithm several times using different random $P$ of order $m$. More than half of the $P\in E[m]$ of order $m$ should generate $E[m]$ as an $\mathcal{O}$-module, since $E[m]\cong\mathcal{O}/m\mathcal{O}$ by Theorem 3.3, and $\mathcal{O}$ can fix at most two order $m$ subgroups. Any such $P$ has self-pairing of order $m$ (by non-degeneracy), so repeating sufficiently often and taking the majority rule answer, this will succeed with overwhelming probability in polynomial time.
@@
Given any basis for $E[m]$, the pairing $\widehat{T}_{m}^{\tau}$ allows us to compute the ‘eigenspaces’, i.e. a basis $P,Q$ such that $[\tau]P\in\langle P\rangle$ and $[\tau]Q\in\langle Q\rangle$. That is, knowing the pairing values on the original basis, we can solve for points with trivial self-pairing.
@@
Consider the elliptic curve $y^{2}=x^{3}+x$ over $\mathbb{F}_{p}$, $p=541$. A basis for $E[5]$ is $P=(109,208)$, $Q=(53,195)$. If we compute the self-pairings $\widehat{T}_{5}^{[i]}([a]P+[b]Q,[a]P+[b]Q)$, for $a,b=0,\ldots,4$, we obtain the following: the left matrix shows the real parts and the right matrix the imaginary parts, taken log base a generator of $\mathbb{F}_{p}$, namely, $48$. So, for example, the second entry on the fourth row indicates that $\widehat{T}_{5}^{[i]}([4]P+[2]Q,[4]P+[2]Q)=(g^{3},g^{4})$.
@@
From this we observe that $\widehat{T}_{5}^{[i]}(P,P)=\widehat{T}_{5}^{[i]}([2]P+Q,[2]P+Q)=1$. This is as dictated by Theorem 4.3, because $[i]P=[3]P$ and $[i]([2]P+Q)=[2]([2]P+Q)$.

### v2

@@ 2.5 Computational assumptions
With regards to computations, we use the word efficient to mean polynomial time in the size of the input, which is itself typically a polynomial in $\log m$ (the torsion) and $\log q$ (where $q$ is the cardinality of the field of definition of the $m$-torsion), in our context. Throughout the paper, when we assume that we are given an $\mathcal{O}$-oriented elliptic curve, we mean that we are given an explicit orientation, and in particular that, given an element $\alpha\in\mathcal{O}$, we can compute its action $[\alpha]$ on a point $P$ on $E$ efficiently.
@@
We assume that $m$ is coprime to the characteristic $p$ of the given field $\mathbb{F}$, and that $m$ is smooth, meaning that its factors are polynomial in size, so that discrete logarithms in $\mu_{m}$ or $E[m]$ are computable in polynomial time. In particular, we can efficiently write any element of $E[m]$ in terms of a given basis.
@@
Let $m>1$ be an integer. Let $E$ be an elliptic curve defined over a field $\mathbb{F}$ (assumed finite in this paper). Suppose that $P\in E(\mathbb{F})[m]$. Choose divisors $D_{P}$ and $D_{Q}$ of disjoint support such that $D_{P}\sim(P)-(\mathcal{O})$ and $D_{Q}\sim(Q)-(\mathcal{O})$. Then $mD_{P}\sim\emptyset$, hence there is a function $f_{P}$ such that $\operatorname{div}(f_{P})=mD_{P}$. The Tate-Lichtenbaum pairing
@@
The standard properties of the Tate pairing are as follows. Proofs can be found in many places, for example [Rob23a] and [CHM+23, Sec 3.2].
@@ 3 Structure of $E[\alpha]$
Suppose first that $\alpha$ is separable. Let $\iota$ be an $\mathcal{O}$-orientation for $E$ and $\mathcal{O^{\prime}}$ be the order of $K$ for which $\iota$ is a primitive orientation. Then as an abelian group the quotient $\operatorname{End}(E)/\mathcal{O^{\prime}}$ is torsion-free and Theorem 3.1 tells us that $E[\alpha]\cong\mathcal{O^{\prime}}/\alpha\mathcal{O^{\prime}}$ as $\mathcal{O}^{\prime}$-modules. We have $N(\alpha)=N(\alpha\mathcal{O^{\prime}})=\left|\mathcal{O^{\prime}}/\alpha\mathcal{O^{\prime}}\right|$, so since $N(\alpha)$ is coprime to $f$, it follows from [Cox22, Proposition 7.18, 7.20] that the natural injection $\mathcal{O}/\alpha\mathcal{O}\rightarrow\mathcal{O^{\prime}}/\alpha\mathcal{O^{\prime}}$ is an isomorphism of $\mathcal{O}$-modules.
@@
Finally, suppose $\alpha$ factors through $[n]$ for some $n>1$ with $n\mid f$. Then as a $\mathbb{Z}$-module, $E[\alpha]\cong\mathbb{Z}/b\mathbb{Z}\times\mathbb{Z}/c\mathbb{Z}$ with $n\mid b\mid c$. Let $Q$ be an arbitrary point of order $c$ in $E[\alpha]$ and extend to a generating set $\{P,Q\}$ for $E[\alpha]$ with $\operatorname{ord}(P)=b,\operatorname{ord}(Q)=c$. Let $\mathcal{O^{\prime}}=\mathbb{Z}[\sigma]$ for some $\sigma$. Then $\mathcal{O}=\mathbb{Z}[f\sigma]$. Since any element of $\mathcal{O}$ is a $\mathbb{Z}$-linear combination of $[1]$ and $[f\sigma]$, whether or not $E[\alpha]$ is cyclic as an $\mathcal{O}$-module is determined by the action of $f\sigma$. We have $[f\sigma]Q=[n\sigma]Q^{\prime}$, where $Q^{\prime}=[f/n]Q$.
If $[\sigma]Q^{\prime}=[s]P+[t]Q$, then $[f\sigma]Q=[ns]P+[nt]Q$, and we cannot obtain $P$ from the action of any $\mathbb{Z}$-linear combination of $[1]$ and $[f\sigma]$ on $Q$ (since $[n]$ is not injective on $E[\alpha]$). Thus, $Q$ cannot be a generator for $E[\alpha]$ as an $\mathcal{O}$-module. Since $Q$ was an arbitrary order $c$ point, and since no point of order strictly less than $c$ can generate $E[\alpha]$ as an $\mathcal{O}$-module (endomorphisms send points of $E$ of order $m$ to points of $E$ of order dividing $m$), $E[\alpha]$ cannot be a cyclic $\mathcal{O}$-module. ∎
@@
Consider the ordinary curve $y^{2}=x^{3}+30x+2$ over $\mathbb{F}_{101}$. Denoting the Frobenius endomorphism of degree $p$ by $\pi$, $\mathbb{Z}[\pi]$ has conductor 2 in the maximal order and $[\mathbb{Z}[\pi]:\mathbb{Z}[\pi^{2}]]=18$. Thus, Theorem 3.3 implies $E[3]$ is not cyclic as a $\mathbb{Z}[\pi^{2}]$-module. Indeed, making a base change to $\mathbb{F}_{101^{2}}$, the 3-torsion of $E$ becomes rational. On the other hand, $E[3]$ is cyclic as a $\mathbb{Z}[\pi]$-module. With $\mathbb{F}_{101^{2}}=\mathbb{F}_{101}(a)$ and $x^{2}-4x+2$ the minimal polynomial of $a$, we have $P=(41a+16,39a+19)\in E[3]$ and $\pi(P)=(60a+79,62a+74)\not\in\langle P\rangle$, hence $\mathbb{Z}[\pi]P=E[3]$.
@@ 4 Sesquilinear pairings
We follow [Sta24] in this section. Suppose $\mathcal{O}=\mathbb{Z}[\tau]$ is an imaginary quadratic order. Let $E$ have CM by $\mathcal{O}$. Let $\rho:\mathcal{O}\rightarrow M_{2\times 2}(\mathbb{Z})$ be the left-regular representation of $\mathcal{O}$ acting on the basis $1$ and $\tau$, i.e.
$\rho(\alpha)=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\quad\iff\quad\alpha=a+c\tau,\quad\alpha\tau=b+d\tau.$
Then we endow the Cartesian square $(\mathbb{F}^{*})^{\times 2}$ of the multiplicative $\mathbb{Z}$-module $\mathbb{F}^{*}$ (i.e. $\mathbb{Z}$-coefficients in the exponent) with a multiplicative $\mathcal{O}$-module action (i.e. $\mathcal{O}$-coefficients in the exponent) via
$(x,y)^{\alpha}=\rho(\alpha)\cdot(x,y)=(x^{a}y^{b},x^{c}y^{d}),\quad\text{where}\quad\rho(\alpha)=\begin{pmatrix}a&b\\ c&d\end{pmatrix}.$ (2)
@@
$\operatorname{div}(f_{P,1})$ $=a([-\tau]P)+b(P)-(a+b)(\infty),$
$\operatorname{div}(f_{P,2})$ $=c([-\tau]P)+d(P)-(c+d)(\infty).$
Choose an auxiliary point $R\in E(\overline{\mathbb{F}})$ and define for $Q\in E(\mathbb{F})$,
@@
Then, choosing $R$ so that the necessary supports are disjoint (i.e. the support of $\operatorname{div}(f_{P,i})$ and $D_{Q,j}$ are disjoint for each pair $i$, $j$), the pairing is defined (using (2)) as
@@
In [Sta24], it is shown how it is possible to think of these definitions as elements of $\mathcal{O}\otimes_{\mathbb{Z}}\operatorname{Pic}^{0}(E)$:
$D_{Q}=D_{Q,1}+\tau\cdot D_{Q,2},\quad D_{P}=([-\tau]P)-(\infty)+\tau\cdot((P)-(\infty)),$
and analogously define $f_{P}$ satisfying $\operatorname{div}(f_{P})={\alpha}\cdot D_{P}$, so that the definition above has the form $f_{P}(D_{Q})$ as for the classical Tate pairing, and the apparent dependence on the basis $1$, $\tau$ disappears. For simplicity here, we stick to the direct definition above. In that same paper, analogous constructions are also given for quaternion orders and Weil-like pairings.
Theorem 4.1 ([Sta24, Theorems 5.4, 5.5, 5.6])
@@
    $\widehat{T}^{\tau}_{N(\alpha)}(P,Q)=\widehat{T}^{\tau}_{\alpha}(P,Q)^{\overline{\alpha}}\pmod{((\mathbb{F}^{*})^{\times 2})^{\alpha}}$
@@
Provided the action of $\tau$ is efficiently computable, then the pairing $\widehat{T}_{\alpha}^{\tau}(P,Q)$ is efficiently computable. That is, it takes polynomially many operations in the field of definition of $P$ and $Q$.
@@
This follows from the definition given above, which is amenable to a Miller-style pairing algorithm; details are in [Sta24, Algorithm 5.7].∎
@@
Consider the image $N(\mathcal{O})$ of $\mathcal{O}$ under the norm map. Then $N(\mathcal{O})$ modulo $m>2$ is a subset of $\{x^{2}:x\in\mathbb{Z}/m\mathbb{Z}\}$ only if $m$ and $\Delta_{\mathcal{O}}$ share a non-trivial factor.
@@
We may assume, by Sunzi’s Theorem (Chinese Remainder Theorem), that $m$ is a prime power $p^{k}>2$. If $p$ is split, the statement follows from the fact that the norm map from $\mathbb{Q}_{p}\otimes_{\mathbb{Z}}\mathcal{O}$ to $\mathbb{Q}_{p}$ is surjective. If $p$ is inert, the norm map is surjective on the residue field, so $N(\mathcal{O})$ modulo $p^{k}$ does include non-squares. ∎
@@
Let $E$ be an elliptic curve oriented by $\mathcal{O}=\mathbb{Z}[\tau]$. Let $m$ be coprime to the discriminant $\Delta_{\mathcal{O}}$. Let $\mathbb{F}$ be a finite field containing the $m$-th roots of unity. Suppose $E[m]=E[m](\mathbb{F})$. Let $P$ have order $m$. Let $s$ be the maximal divisor of $m$ such that $E[s]\subseteq\mathcal{O}P$. Then the multiplicative order $m^{\prime}$ of $\widehat{T}_{m}^{\tau}(P,P)$ satisfies $s\mid m^{\prime}\mid 2s^{2}$. In particular, if $\mathcal{O}P=E[m]$, then $s=m$ and the self-pairing has order $m$. If $\mathcal{O}P=\mathbb{Z}P$, then $s=1$, and in fact, in this case, the self-pairing is trivial.
To rephrase the last sentence, the self-pairing is trivial on the eigenspaces for the action of $\mathcal{O}$ on $E[m]$. This observation by itself is a consequence of the classification of self-pairings in [CHM+23].
@@
Let $m^{\prime}\mid m$ be the order of $\widehat{T}_{m}^{\tau}(P,P)$. Suppose $s$ is the maximal divisor of $m$ so that $E[s]\subseteq\mathcal{O}P$. In other words, $\mathcal{O}P\cong\mathbb{Z}/s\mathbb{Z}\times\mathbb{Z}/m\mathbb{Z}$ and $\mathcal{O}P/\mathbb{Z}P\cong\mathbb{Z}/s\mathbb{Z}$ as abelian groups. In particular, $[s]\mathcal{O}P\in\mathbb{Z}P$. Thus $\mathcal{O}[s]P=\mathbb{Z}[s]P$.
We will show that $m^{\prime}\mid 2s^{2}$ and $s\mid m^{\prime}$. Let $\lambda\in\mathcal{O}$. Then $[\lambda s]P=[ks]P$ for some $k=k(\lambda)\in\mathbb{Z}$, and then
@@
Ranging over all $\lambda\in\mathcal{O}$, we conclude that $N(\lambda)$ are squares modulo $m^{\prime\prime}:=m^{\prime}/\gcd(m^{\prime},s^{2})$, the multiplicative order of $\widehat{T}_{m}^{\tau}([s]P,[s]P)$, contradicting that $m$ is coprime to the discriminant unless $m^{\prime\prime}=1$ or $2$ by Lemma 1. Therefore $m^{\prime}\mid 2s^{2}$. In the case where $s=1$, this argument implies only that the order of $\widehat{T}^{\tau}_{m}(P,P)$ is at most 2. However, the fact that in this case the order of the self-pairing is trivial follows immediately from [CHM+23, Proposition 4.8].
@@
This has order $s$ on the left. Therefore $\widehat{T}_{m}^{\tau}(P,P)$ must have order a multiple of $s$. Hence $s\mid m^{\prime}$. ∎
@@
As discussed in Section 2.3, the authors of [CHM+23], the authors show that non-trivial cyclic self-pairings can only exist for $P$ of order $m$ dividing $\Delta_{\mathcal{O}}$. The reason our pairings are not ruled out by this result is that our pairings are defined not only on cyclic subgroups stabilized by the orientation (where they are in fact trivial, as required).
The following is a partial converse to Theorem 4.2.
@@
Let $E$ be an elliptic curve defined over a finite field $\mathbb{F}$, and let $m\in\mathbb{Z}$. Let a basis for $E[m]$ be given. Suppose arithmetic in $\mathbb{F}$, discrete logarithms in $\mathbb{F}^{*}$ modulo $m$, and group law computations on $E[m]$ can all be accomplished in polynomial time. Suppose $\varphi(m)>\sqrt{2/3}m$. Suppose $E$ is known to be oriented by $\mathcal{O}=\mathbb{Z}[\tau]$ (but the orientation $\iota$ is not given), and suppose $m$ is coprime to the discriminant $\Delta_{\mathcal{O}}$. Then the computation of arbitrary pairings $\widehat{T}^{\tau}_{m}(P,Q)$ on $E[m]$ is Monte-Carlo equivalent in polynomial time to the computation of the action of $[\tau]$ on $E[m]$.
By Monte-Carlo equivalent, we mean that there is an arbitrarily small probability that the algorithm will return incorrectly. The condition on $\varphi(m)$ can be improved: what should be required is that $\varphi(m)$ non-negligibly exceed $m/\sqrt{2}$.
@@
Note that computation of $[\tau]$ allows for computation of $[\overline{\tau}]=[Tr(\tau)]-[\tau]$.
@@
Conversely, suppose one can compute $\widehat{T}^{\tau}_{m}(P,Q)$ for any $P,Q\in E[m]$. We will show how to compute the action of $[\tau]$ on $E[m]$. The pairing is non-degenerate as a consequence of the given hypotheses. It is possible to sample randomly from the subset of order $m$ points in $E[m]$, by choosing $P$ uniformly randomly as a linear combination $aP_{1}+bP_{2}$ of the given basis $P_{1},P_{2}$ such that $\gcd(a,b)$ is coprime to $m$. Choose such a $P$ of order $m$ and compute $\widehat{T}^{\tau}_{m}(P,P)$.
@@
If $\mathcal{O}P\neq E[m]$, then the algorithm is not guaranteed to be correct. Therefore, we run the algorithm several times using different random $P$ of order $m$. We have $E[m]\cong\mathcal{O}/m\mathcal{O}$ by Theorem 3.3. Any element of $\mathcal{O}$ is a $\mathcal{O}$-module generator of $\mathcal{O}/m\mathcal{O}$ provided it is coprime to $m$ (since $1$ is a generator and it has an inverse modulo $m$). So the proportion of such generators is at least $(\varphi(m)/m)^{2}$. By our assumption on $m$, this exceeds $2/3$. Any such $P$ has self-pairing of order $m$ (by non-degeneracy), so repeating sufficiently often and taking the majority rule answer, this will succeed with overwhelming probability in polynomial time. ∎
@@
If $\varphi(m)/m$ is non-negligible, then one can sample points uniformly at random and use the pairing to check whether they generate $\mathcal{O}$, with a high probability of success. However, if $m$ is badly behaved, for example, a primorial, then $\varphi(m)/m$ may be less than $1/m^{x}$ for some $x>0$.
Remark 4
Given any basis for $E[m]$, the pairing $\widehat{T}_{m}^{\tau}$ allows us to compute the ‘eigenspaces’, i.e. a basis $P,Q$ such that $[\tau]P\in\mathbb{Z}P$ and $[\tau]Q\in\mathbb{Z}Q$. That is, knowing the pairing values on the original basis, we can solve for points with trivial self-pairing.
@@
Consider the elliptic curve $y^{2}=x^{3}+x$ over $\mathbb{F}_{p}$, $p=541$. A basis for $E[5]$ is $P=(109,208)$, $Q=(53,195)$. If we compute the self-pairings $\widehat{T}_{5}^{[i]}([a]P+[b]Q,[a]P+[b]Q)$, for $a,b=0,\ldots,4$, we obtain the following: the left matrix shows the real parts and the right matrix the imaginary parts, taken to the $\log$ base 48 (48 is a generator of $\mathbb{F}_{541}^{*}$). So, for example, the fourth entry ($a=3$) in the second column ($b=1$) (of both matrices) indicates that $\widehat{T}_{5}^{[i]}([3]P+[1]Q,[3]P+[1]Q)=(g^{3},g^{4})$.
@@
We can also read off, for example, that $\widehat{T}_{5}^{[i]}(P,P)=\widehat{T}_{5}^{[i]}([2]P+Q,[2]P+Q)=1$. Thus the matrices have zeroes on the first column and on the coordinates $(a,b)$ which are multiples of $(2,1)$ modulo $5$. This is as dictated by Theorem 4.3, because $P\in E[2+i]$ and $[2]P+Q\in E[2-i]$, which implies $[i]P=[3]P$ and $[i]([2]P+Q)=[2]([2]P+Q)$. In other words, the subgroups $E[2\pm i]\cong\mathcal{O}/(2\pm i)\mathcal{O}$ are the eigenspaces for the action of $\mathcal{O}$ on $E[5]\cong\mathcal{O}/5\mathcal{O}$.